# Best Practices for Biorisk Evaluations on Open-Weight Bio-Foundation Models

### v3

@@ 1 Introduction
The growing capabilities of bio-foundation models have raised concerns about their potential misuse [43, 49, 52, 13]. This concern is particularly prominent for open-weight models, which allow greater freedom for adversarial modifications, especially when fine-tuned for malicious purposes [45, 44].
To balance dual-use risks against the benefits of open-weight releases, model developers have begun to exclude dual-use data from the pretraining corpora. For instance, the OpenGenome dataset [^1] https://huggingface.co/datasets/arcinstitute/opengenome2, used to train the Evo model family, intentionally excludes eukaryotic viral sequences [37, 5]. As a result, Evo models exhibit poor initial performance in predicting mutational effects on such sequences [5].
Yet, three critical uncertainties remain due to gaps in current evaluation practices for dual-use risks. First, while some initial efforts have been made in examining biorisks in natural language models [19, 56, 31], there is no comparable evaluation framework for bio-foundation models, making systematic dual-use risk assessment challenging. Second, the effectiveness of data filtering for bio-foundation models has not been fully assessed under a threat model where adversaries can fine-tune the model. Data filtering has sometimes been a robust approach for reducing dual-use risks in natural language models, even after some adversarial fine-tuning [39, 7]. But to date, there has not been a similar assessment in the bio-foundation model context to determine whether the filtered capabilities are trivial to re-learn. Third, elicitation practices for testing the safety of open-weight bio-foundation models are underexplored. Simple methods including probing have not yet been tried on these models, leaving open the possibility that latent representations still encode the necessary knowledge to enable misuse [31, 22].
@@ 2 Related Work
Deployment  [59] ? (Afsnit˜4.2, 4.3)
@@
Bio-foundation models are large models trained on diverse biological datasets to learn generalized representations that can be adapted to various downstream tasks [18]. They can be categorized by the type of biomolecular data they process, including Genomic Models (models that operate on genomic sequences, e.g., Enformer [3], DNABert [60, 23], Evo family [37, 5]), Transcriptomics Models (models that analyze high-dimensional gene expression data, e.g., Geneformer [8], CellFM [58]), and Protein Models (models that handle protein sequences, e.g., Alphafold [26], ESM 2 [33], AMPLIFY [16]). In this paper, we focus on Genomic foundation models due to their general strong capability and wide application domains. As illustrated in Tabel˜1, bio-foundation models are used not only for sequence generation but also for regression and classification tasks. These include applications such as mutational effect prediction, clinical variant interpretation, and related functional genomics analyses.
@@
Data filtering has been widely adopted during the pretraining stage of open-weight foundation models to mitigate legal and safety risks. In natural language models, it has proven effective in reducing harmful outputs [2], limiting private information leakage [27], and mitigating copyright issues [36]. Recent studies have further advanced filtering methods by applying harmfulness classifiers [7]. Compared to post-training safeguards such as circuit breakers [61], data filtering has emerged as a more robust approach in language models [39]. In the domain of bio-foundation models, data filtering is likewise employed to reduce misuse risks. For example, Evo 2 excluded eukaryotic viral data during pretraining [37, 5] to mitigate potential biosecurity concerns.
@@
Various efforts have been made to explore dual-use risks in bio-foundation models. For example, Evo 2 assesses such risks during the deployment stage using metrics like perplexity distribution, mutational effect prediction, protein generation success rates, and ancestry bias in eukaryotic viral sequences [5]; Genebreaker [59] investigates the ability to induce bio-foundation models to generate eukaryotic viral sequences via inference-time guided search. Beyond bio-foundation models, several studies have examined dual-use concerns in Language Foundation Models. For instance, system cards from OpenAI model families [41, 42] assess the biological dual-use risk through evaluations on long-form bio-risk questions, multimodal virology troubleshooting, ProtocolQA [29], and tacit knowledge probing. In terms of malicious fine-tuning risks, Wallace et al. [51] proposes a worst-case estimation approach that stress-tests maximum knowledge gains after fine-tuning with adequate compute budgets. However, to the best of our knowledge, there is a lack of research assessing the dual-use risks associated with fine-tuning bio-foundation models – a gap this paper aims to address.
@@ 3.1 Threat Model
Following the insights from the prior work [5, 57], we consider a threat model where adversaries seek to exploit open-weight bio-foundation models to design pathogenic eukaryotic viruses. From the adversaries’ perspective, the objective is to leverage the model to facilitate pathogen engineering. Specifically, adversaries may (1) generate novel and viable viral sequences; (2) perform targeted in silico optimization to enhance hazardous traits such as protein stability or receptor binding affinity, and (3) rank and select potential candidates based on the predicted virulence. Full access to model weights further enables adversaries to fine-tune the model on public biological datasets or probe hidden representations to improve performance on these tasks. From the defender’s perspective, the objective is to minimize the risk of such misuse after releasing the model weights. To this end, pre-release safety approaches will be applied to constrain malicious utility while maintaining scientific value for legitimate users. Our safety evaluations are thus designed to assess the model’s harmful capabilities across the adversarial tasks outlined above.
@@ 3.2 BioRiskEval
We evaluate mutational effect prediction using 16 Deep Mutational Scanning (DMS) datasets [17] on human-infecting viruses from ProteinGym [38]. Each DMS dataset consists of mutant variants with experimentally measured fitness scores obtained through high-throughput selection assays (e.g., growth, binding, or expression). Fitness scores are estimated by comparing pre- and post-selection sequencing counts, normalized to the wild type. We assess models by computing the absolute Spearman rank correlation $|\rho|$ between the experimental fitness values and the model-derived sequence scores for the same mutants. A larger $|\rho|$ indicates better agreement with experimental rankings and thus a stronger ability to capture mutational effects. We convert ProteinGym’s protein sequences into nucleotide sequences for evaluations with Evo 2, in a process detailed in afsnit˜A.2. More details on the DMS datasets can be found in tabel˜4.
@@ 4 Experimental Results
We use Evo2-7B model [5], a genome foundation model designed for both nucleotide sequence modeling and biological representation prediction, as our primary model due to its strong performance across diverse downstream tasks. Moreover, its deliberate exclusion of eukaryotic viral sequences during pre-training enables us to assess the effectiveness of data filtering, offering a controlled safety baseline for our experiments. See Appendiks˜B for more details.
@@ 4.1.1 Inter-Species Generalizability
We plot the perplexity distribution across various fine-tuning steps in Figur˜2(a). Notably, for both case studies, the perplexity distribution on the hold-out species quickly dropped to the same level as on IMG/PR within 50 fine-tuning steps, which is merely 0.72 H100 GPU hours under our experiment settings. This indicates that the model can easily generalize to the other species within the same genus after a few steps of fine-tuning. Given the typically high structural similarity across species in the same genus [4, 50, 40], such a result is not surprising. However, it highlights that data filtering is not tamper-resistant in this setting: excluding one species does not robustly prevent efficient recovery of capabilities through fine-tuning on related species. Moreover, this also implies that malicious fine-tuning could be significantly streamlined, requiring only a subset of representative species rather than exhaustive coverage.
@@ 4.2.1 Log-Likelihood-Based Mutational Effect Prediction
Setup. Following prior work [35, 38], we score mutational effects with Evo2-7B model by calculating log-likelihood $\mathcal{S}_{LL}=\sum_{j=1}^{L}\log P(x_{j}^{mt}|x_{<j}^{mt},\theta)$ for each mutation $x^{mt}$. Model predictions are compared against experimental DMS scores using Spearman correlation $\rho$, where higher absolute correlation indicates stronger agreement with measured mutational predictive power. To investigate the improvement through fine-tuning, we fine-tune the model on the longest sequences for each species in BioRiskEval-Gen for 100 to 2,000 steps (See Appendiks˜A for more details). As baseline, we also include zero-shot scoring with ESM 2, a BERT-style protein language model trained without filtering eukaryotic viral sequences [33]. For ESM 2, we adopt masked marginal scoring, which is defined as $\mathcal{S}_{MM}=\sum_{i\in M}\log p(x_{i}=x_{i}^{mt}|x_{-i}^{mt})-\log p(x_{i}=x_{i}^{wt}|x_{-i}^{wt})$, where $M$ is the set of mutation positions, $x_{-i}$ is the sequence with place $i$ masked, $x^{mt}$ is the mutation sequence, and $x^{wt}$ is the wildtype sequence.
@@ 4.2.2 Predicting Mutational Effect with Linear Probe
Setup. To test whether hidden representations encode knowledge relevant for predicting mutational effects, we perform linear probing on model hidden states to predict continuous DMS scores. From BioRiskEval-Mut, we apply stratified sampling based on the DMS score, and sample 500 mutations from each of the 14 DMS datasets (excluding two with fewer than 500 mutations), allocating 400 mutations per dataset for training and 100 for validation, while all remaining mutations are used as the test set. This yields 5,600 training, 1,400 validation, and 148,505 test mutations (denoted as BioRiskEval-Mut-Probe, see Tabel˜5 for more details), with probes fit on only 3.6% of the DMS data and evaluated on 95.4%. Here, the probes are universal and are trained across all pooled datasets using the closed-form solution with a mean square error objective. Similar to Afsnit˜4.2.1, the performance is measured by the absolute Spearman correlation $|\rho|$. Since fine-tuning will alter the hidden representations, we do not predefine a probe layer; instead, for each checkpoint, we train probes on all layers and select (i) the layer with the lowest training root mean square error (RMSE) and (ii) the highest validation Spearman correlation, and report the corresponding test performance in terms of $|\rho|$. For comparison, we not only compute log-likelihood-based mean $|\rho|$ on the test set of BioRiskEval-Mut-Probe for Evo2-7B and ESM 2 checkpoints, but also probe the hidden-layer feature from LLaMA-3.1-8B-Instruct [11], a natural language model that is not trained on the nucleotide sequences, and shares the same number of layers and hidden layer dimension with Evo2-7B. By comparing the performance between Evo2-7B and LLaMA-3.1-8B-Instruct, we will know how much performance uplift adversaries can achieve with the assistance of bio-foundation models.
@@ Conclusion.
In this study, we introduce BioRiskEval framework, which offers a comprehensive assessment of the dual-use risks in bio-foundation models across three dimensions. Our experiments with Evo 2 reveal that data filtering is not tamper-resistant under certain circumstances, and may fail to prevent the model from learning malicious capabilities like mutational effect prediction, virulence prediction in the latent space during the pre-training stage. That said, our intention is not to assert that data filtering is completely ineffective, but to highlight the scenarios where it may fall short—underscoring the risks of relying on it as the sole defense mechanism. These findings call for more robust safety strategies for open-weight bio-foundation models that go beyond data filtering alone. We further argue that future safety to open-weight bio-foundation models should account for adversarial manipulations, such as fine-tuning and probing, which are practical for adversaries and can potentially expand the “bubble of risk” [53].
@@ A.2 Protein Sequences to Nucleotide Sequences Conversion Pipeline
We first find wild type nucleotide sequence by using the NBCI BLAST program in three ways: exact match with organism filter, exact match, and 99% match with seeded random codon replacement for the rest of the amino acids [1]. After the wild type is found or constructed, we perform the mutation swap as detailed by the original DMS protein dataset, and seed the random selection of codon for each replacement at a time.
Our pipeline do not recreate exact nucleotide sequences that Evo 2 may evaluated on due to the random picking among viable codons, and 2 DMS files that we did not reconstruct due to BLAST not finding wild type nucleotide with high match. We consider this a valid approach because DMS experiments artificially introduce mutations biologically. Furthermore, experimentally we reproduce the mean absolute Spearman rank for human viruses on Evo2-7B, with less than 0.5 mean absolute Spearman rank matching S2B in [5].
@@ A.4 DMS Datset Overview
A0A192B1T2_9HIV1_Haddox_2018 [20] 12,577 2,556 HIV-1 (isolate BF520.W14M.C2)
A0A2Z5U3Z0_9INFA_Doud_2016 [9] 10,715 1,695 Influenza A virus (A/WSN/1933(H1N1))
A0A2Z5U3Z0_9INFA_Wu_2014 [54] 2,350 1,695 Influenza A virus (A/WSN/1933(H1N1))
A4D664_9INFA_Soh_2019 [47] 14,421 2,277 Influenza A virus (A/green-winged teal/Ohio/175/1986(H2N1))
C6KNH7_9INFA_Lee_2018 [30] 10,754 1,698 Influenza A virus (A/Perth/16/2009(H3N2))
CAPSD_AAV2S_Sinai_2021 [46] 42,328 2,205 Adeno-associated virus 2 (AAV-2) (isolate Srivastava/1982)
ENV_HV1B9_DuenasDecamp_2016 [12] 375 2,559 HIV-1 group M subtype B (strain 89.6)
I6TAH8_I68A0_Doud_2015 [10] 9,462 1,494 Influenza A virus (A/Aichi/2/1968 H3N2)
NRAM_I33A0_Jiang_2016 [24] 298 1,359 Influenza A virus (A/Wilson-Smith/1933 H1N1)
PA_I34A1_Wu_2015 [55] 1,820 2,148 Influenza A virus (A/Puerto Rico/8/1934 H1N1)
POLG_DEN26_Suphatrakul_2023 [48] 16,897 2,700 Dengue virus 2 (DENV-2) (strain Thailand/16681/1984)
Q2N0S5_9HIV1_Haddox_2018 [20] 12,729 2,580 HIV-1 (isolate BG505.W6M.C2.T332N)
R1AB_SARS2_Flynn_2022 [15] 5,725 918 SARS-CoV-2 (isolate MN908947.3, RefSeq NC_045512.2)
RDRP_I33A0_Li_2023 [32] 12,003 2,271 Influenza A virus (A/Wilson-Smith/1933 H1N1)
REV_HV1H2_Fernandes_2016 [14] 2,147 348 HIV-1 group M subtype B (isolate HXB2)
TAT_HV1BR_Fernandes_2016 [14] 1,577 258 HIV-1 group M subtype B (isolate BRU/LAI)
@@
A0A192B1T2_9HIV1_Haddox_2018 [20] 400 100 12,077 12,577
A0A2Z5U3Z0_9INFA_Doud_2016 [9] 400 100 10,215 10,715
A0A2Z5U3Z0_9INFA_Wu_2014 [54] 400 100 1,850 2,350
A4D664_9INFA_Soh_2019 [47] 400 100 13,921 14,421
C6KNH7_9INFA_Lee_2018 [30] 400 100 10,254 10,754
CAPSD_AAV2S_Sinai_2021 [46] 400 100 41,828 42,328
I6TAH8_I68A0_Doud_2015 [10] 400 100 8,962 9,462
PA_I34A1_Wu_2015 [55] 400 100 1,320 1,820
POLG_DEN26_Suphatrakul_2023 [48] 400 100 16,397 16,897
Q2N0S5_9HIV1_Haddox_2018 [20] 400 100 12,229 12,729
R1AB_SARS2_Flynn_2022 [15] 400 100 5,225 5,725
RDRP_I33A0_Li_2023 [32] 400 100 11,503 12,003
REV_HV1H2_Fernandes_2016 [14] 400 100 1,647 2,147
TAT_HV1BR_Fernandes_2016 [14] 400 100 1,077 1,577

### v4

@@ 1 Introduction
The growing capabilities of bio-foundation models have raised concerns about their potential misuse [puzis2020increased, urbina2022dual, wang2025call, feldman2025resilient]. This concern is particularly prominent for open-weight models, which allow greater freedom for adversarial modifications, especially when fine-tuned for malicious purposes [qifine, qi2024evaluating].
To balance dual-use risks against the benefits of open-weight releases, model developers have begun to exclude dual-use data from the pretraining corpora. For instance, the OpenGenome dataset [^1] https://huggingface.co/datasets/arcinstitute/opengenome2, used to train the Evo model family, intentionally excludes eukaryotic viral sequences [nguyen2024sequence, brixi2025genome]. As a result, Evo models exhibit poor initial performance in predicting mutational effects on such sequences [brixi2025genome].
Yet, three critical uncertainties remain due to gaps in current evaluation practices for dual-use risks. First, while some initial efforts have been made in examining biorisks in natural language models [gotting2025virology, yin2025genome, li2024wmdp], there is no comparable evaluation framework for bio-foundation models, making systematic dual-use risk assessment challenging. Second, the effectiveness of data filtering for bio-foundation models has not been fully assessed under a threat model where adversaries can fine-tune the model. Data filtering has sometimes been a robust approach for reducing dual-use risks in natural language models, even after some adversarial fine-tuning [obrien2025deepignorancefilteringpretraining, chen2025enhancing]. But to date, there has not been a similar assessment in the bio-foundation model context to determine whether the filtered capabilities are trivial to re-learn. Third, elicitation practices for testing the safety of open-weight bio-foundation models are underexplored. Simple methods including probing have not yet been tried on these models, leaving open the possibility that latent representations still encode the necessary knowledge to enable misuse [li2024wmdp, jang2025prompt].
@@ 2 Related Work
Deployment  [zhang2025genebreaker] ? (Afsnit˜4.2, 4.3)
@@
Bio-foundation models are large models trained on diverse biological datasets to learn generalized representations that can be adapted to various downstream tasks [fu2025foundation]. They can be categorized by the type of biomolecular data they process, including Genomic Models (models that operate on genomic sequences, e.g., Enformer [avsec2021effective], DNABert [zhou2023dnabert, ji2021dnabert], Evo family [nguyen2024sequence, brixi2025genome]), Transcriptomics Models (models that analyze high-dimensional gene expression data, e.g., Geneformer [cui2024geneformer], CellFM [zeng2025cellfm]), and Protein Models (models that handle protein sequences, e.g., Alphafold [jumper2021highly], ESM 2 [lin2023evolutionary], AMPLIFY [fournier2024protein]). In this paper, we focus on Genomic foundation models due to their general strong capability and wide application domains. As illustrated in Tabel˜1, bio-foundation models are used not only for sequence generation but also for regression and classification tasks. These include applications such as mutational effect prediction, clinical variant interpretation, and related functional genomics analyses.
@@
Data filtering has been widely adopted during the pretraining stage of open-weight foundation models to mitigate legal and safety risks. In natural language models, it has proven effective in reducing harmful outputs [anil2023palm], limiting private information leakage [korbak2023pretraining], and mitigating copyright issues [min2023silo]. Recent studies have further advanced filtering methods by applying harmfulness classifiers [chen2025enhancing]. Compared to post-training safeguards such as circuit breakers [zou2024improving], data filtering has emerged as a more robust approach in language models [obrien2025deepignorancefilteringpretraining]. In the domain of bio-foundation models, data filtering is likewise employed to reduce misuse risks. For example, Evo 2 excluded eukaryotic viral data during pretraining [nguyen2024sequence, brixi2025genome] to mitigate potential biosecurity concerns.
@@
Various efforts have been made to explore dual-use risks in bio-foundation models. For example, Evo 2 assesses such risks during the deployment stage using metrics like perplexity distribution, mutational effect prediction, protein generation success rates, and ancestry bias in eukaryotic viral sequences [brixi2025genome]; Genebreaker [zhang2025genebreaker] investigates the ability to induce bio-foundation models to generate eukaryotic viral sequences via inference-time guided search. Beyond bio-foundation models, several studies have examined dual-use concerns in Language Foundation Models. For instance, system cards from OpenAI model families [openai2025gpt5, openai2025o3] assess the biological dual-use risk through evaluations on long-form bio-risk questions, multimodal virology troubleshooting, ProtocolQA [laurent2024lab], and tacit knowledge probing. In terms of malicious fine-tuning risks, wallace2025estimating proposes a worst-case estimation approach that stress-tests maximum knowledge gains after fine-tuning with adequate compute budgets. However, to the best of our knowledge, there is a lack of research assessing the dual-use risks associated with fine-tuning bio-foundation models – a gap this paper aims to address.
@@ 3.1 Threat Model
Following the insights from the prior work [brixi2025genome, yin2023vipal], we consider a threat model where adversaries seek to exploit open-weight bio-foundation models to design pathogenic eukaryotic viruses. From the adversaries’ perspective, the objective is to leverage the model to facilitate pathogen engineering. Specifically, adversaries may (1) generate novel and viable viral sequences; (2) perform targeted in silico optimization to enhance hazardous traits such as protein stability or receptor binding affinity, and (3) rank and select potential candidates based on the predicted virulence. Full access to model weights further enables adversaries to fine-tune the model on public biological datasets or probe hidden representations to improve performance on these tasks. From the defender’s perspective, the objective is to minimize the risk of such misuse after releasing the model weights. To this end, pre-release safety approaches will be applied to constrain malicious utility while maintaining scientific value for legitimate users. Our safety evaluations are thus designed to assess the model’s harmful capabilities across the adversarial tasks outlined above.
@@ 3.2 BioRiskEval
We evaluate mutational effect prediction using 16 Deep Mutational Scanning (DMS) datasets [fowler2014deep] on human-infecting viruses from ProteinGym [notin2023proteingym]. Each DMS dataset consists of mutant variants with experimentally measured fitness scores obtained through high-throughput selection assays (e.g., growth, binding, or expression). Fitness scores are estimated by comparing pre- and post-selection sequencing counts, normalized to the wild type. We assess models by computing the absolute Spearman rank correlation $|\rho|$ between the experimental fitness values and the model-derived sequence scores for the same mutants. A larger $|\rho|$ indicates better agreement with experimental rankings and thus a stronger ability to capture mutational effects. We convert ProteinGym’s protein sequences into nucleotide sequences for evaluations with Evo 2, in a process detailed in afsnit˜A.2. More details on the DMS datasets can be found in tabel˜4.
@@ 4 Experimental Results
We use Evo2-7B model [brixi2025genome], a genome foundation model designed for both nucleotide sequence modeling and biological representation prediction, as our primary model due to its strong performance across diverse downstream tasks. Moreover, its deliberate exclusion of eukaryotic viral sequences during pre-training enables us to assess the effectiveness of data filtering, offering a controlled safety baseline for our experiments. See Appendiks˜B for more details.
@@ 4.1.1 Inter-Species Generalizability
We plot the perplexity distribution across various fine-tuning steps in Figur˜2(a). Notably, for both case studies, the perplexity distribution on the hold-out species quickly dropped to the same level as on IMG/PR within 50 fine-tuning steps, which is merely 0.72 H100 GPU hours under our experiment settings. This indicates that the model can easily generalize to the other species within the same genus after a few steps of fine-tuning. Given the typically high structural similarity across species in the same genus [baines2007chapter, van2010evolution, iarc2007human], such a result is not surprising. However, it highlights that data filtering is not tamper-resistant in this setting: excluding one species does not robustly prevent efficient recovery of capabilities through fine-tuning on related species. Moreover, this also implies that malicious fine-tuning could be significantly streamlined, requiring only a subset of representative species rather than exhaustive coverage.
@@ 4.2.1 Log-Likelihood-Based Mutational Effect Prediction
Setup. Following prior work [meier2021language, notin2023proteingym], we score mutational effects with Evo2-7B model by calculating log-likelihood $\mathcal{S}_{LL}=\sum_{j=1}^{L}\log P(x_{j}^{mt}|x_{<j}^{mt},\theta)$ for each mutation $x^{mt}$. Model predictions are compared against experimental DMS scores using Spearman correlation $\rho$, where higher absolute correlation indicates stronger agreement with measured mutational predictive power. To investigate the improvement through fine-tuning, we fine-tune the model on the longest sequences for each species in BioRiskEval-Gen for 100 to 2,000 steps (See Appendiks˜A for more details). As baseline, we also include zero-shot scoring with ESM 2, a BERT-style protein language model trained without filtering eukaryotic viral sequences [lin2023evolutionary]. For ESM 2, we adopt masked marginal scoring, which is defined as $\mathcal{S}_{MM}=\sum_{i\in M}\log p(x_{i}=x_{i}^{mt}|x_{-i}^{mt})-\log p(x_{i}=x_{i}^{wt}|x_{-i}^{wt})$, where $M$ is the set of mutation positions, $x_{-i}$ is the sequence with place $i$ masked, $x^{mt}$ is the mutation sequence, and $x^{wt}$ is the wildtype sequence.
@@ 4.2.2 Predicting Mutational Effect with Linear Probe
Setup. To test whether hidden representations encode knowledge relevant for predicting mutational effects, we perform linear probing on model hidden states to predict continuous DMS scores. From BioRiskEval-Mut, we apply stratified sampling based on the DMS score, and sample 500 mutations from each of the 14 DMS datasets (excluding two with fewer than 500 mutations), allocating 400 mutations per dataset for training and 100 for validation, while all remaining mutations are used as the test set. This yields 5,600 training, 1,400 validation, and 148,505 test mutations (denoted as BioRiskEval-Mut-Probe, see Tabel˜5 for more details), with probes fit on only 3.6% of the DMS data and evaluated on 95.4%. Here, the probes are universal and are trained across all pooled datasets using the closed-form solution with a mean square error objective. Similar to Afsnit˜4.2.1, the performance is measured by the absolute Spearman correlation $|\rho|$. Since fine-tuning will alter the hidden representations, we do not predefine a probe layer; instead, for each checkpoint, we train probes on all layers and select (i) the layer with the lowest training root mean square error (RMSE) and (ii) the highest validation Spearman correlation, and report the corresponding test performance in terms of $|\rho|$. For comparison, we not only compute log-likelihood-based mean $|\rho|$ on the test set of BioRiskEval-Mut-Probe for Evo2-7B and ESM 2 checkpoints, but also probe the hidden-layer feature from LLaMA-3.1-8B-Instruct [dubey2024llama], a natural language model that is not trained on the nucleotide sequences, and shares the same number of layers and hidden layer dimension with Evo2-7B. By comparing the performance between Evo2-7B and LLaMA-3.1-8B-Instruct, we will know how much performance uplift adversaries can achieve with the assistance of bio-foundation models.
@@ Conclusion.
In this study, we introduce BioRiskEval framework, which offers a comprehensive assessment of the dual-use risks in bio-foundation models across three dimensions. Our experiments with Evo 2 reveal that data filtering is not tamper-resistant under certain circumstances, and may fail to prevent the model from learning malicious capabilities like mutational effect prediction, virulence prediction in the latent space during the pre-training stage. That said, our intention is not to assert that data filtering is completely ineffective, but to highlight the scenarios where it may fall short—underscoring the risks of relying on it as the sole defense mechanism. These findings call for more robust safety strategies for open-weight bio-foundation models that go beyond data filtering alone. We further argue that future safety to open-weight bio-foundation models should account for adversarial manipulations, such as fine-tuning and probing, which are practical for adversaries and can potentially expand the “bubble of risk” [wei2025dynamic].
@@ A.2 Protein Sequences to Nucleotide Sequences Conversion Pipeline
We first find wild type nucleotide sequence by using the NBCI BLAST program in three ways: exact match with organism filter, exact match, and 99% match with seeded random codon replacement for the rest of the amino acids [altschul1990basic]. After the wild type is found or constructed, we perform the mutation swap as detailed by the original DMS protein dataset, and seed the random selection of codon for each replacement at a time.
Our pipeline do not recreate exact nucleotide sequences that Evo 2 may evaluated on due to the random picking among viable codons, and 2 DMS files that we did not reconstruct due to BLAST not finding wild type nucleotide with high match. We consider this a valid approach because DMS experiments artificially introduce mutations biologically. Furthermore, experimentally we reproduce the mean absolute Spearman rank for human viruses on Evo2-7B, with less than 0.5 mean absolute Spearman rank matching S2B in [brixi2025genome].
@@ A.4 DMS Datset Overview
A0A192B1T2_9HIV1_Haddox_2018 [haddox2018mapping] 12,577 2,556 HIV-1 (isolate BF520.W14M.C2)
A0A2Z5U3Z0_9INFA_Doud_2016 [doud2016accurate] 10,715 1,695 Influenza A virus (A/WSN/1933(H1N1))
A0A2Z5U3Z0_9INFA_Wu_2014 [wu2014high] 2,350 1,695 Influenza A virus (A/WSN/1933(H1N1))
A4D664_9INFA_Soh_2019 [soh2019comprehensive] 14,421 2,277 Influenza A virus (A/green-winged teal/Ohio/175/1986(H2N1))
C6KNH7_9INFA_Lee_2018 [lee2018deep] 10,754 1,698 Influenza A virus (A/Perth/16/2009(H3N2))
CAPSD_AAV2S_Sinai_2021 [sinai2021generative] 42,328 2,205 Adeno-associated virus 2 (AAV-2) (isolate Srivastava/1982)
ENV_HV1B9_DuenasDecamp_2016 [duenas2016saturation] 375 2,559 HIV-1 group M subtype B (strain 89.6)
I6TAH8_I68A0_Doud_2015 [doud2015site] 9,462 1,494 Influenza A virus (A/Aichi/2/1968 H3N2)
NRAM_I33A0_Jiang_2016 [jiang2016balance] 298 1,359 Influenza A virus (A/Wilson-Smith/1933 H1N1)
PA_I34A1_Wu_2015 [wu2015functional] 1,820 2,148 Influenza A virus (A/Puerto Rico/8/1934 H1N1)
POLG_DEN26_Suphatrakul_2023 [suphatrakul2023functional] 16,897 2,700 Dengue virus 2 (DENV-2) (strain Thailand/16681/1984)
Q2N0S5_9HIV1_Haddox_2018 [haddox2018mapping] 12,729 2,580 HIV-1 (isolate BG505.W6M.C2.T332N)
R1AB_SARS2_Flynn_2022 [flynn2022comprehensive] 5,725 918 SARS-CoV-2 (isolate MN908947.3, RefSeq NC_045512.2)
RDRP_I33A0_Li_2023 [li2023deep] 12,003 2,271 Influenza A virus (A/Wilson-Smith/1933 H1N1)
REV_HV1H2_Fernandes_2016 [fernandes2016functional] 2,147 348 HIV-1 group M subtype B (isolate HXB2)
TAT_HV1BR_Fernandes_2016 [fernandes2016functional] 1,577 258 HIV-1 group M subtype B (isolate BRU/LAI)
@@
A0A192B1T2_9HIV1_Haddox_2018 [haddox2018mapping] 400 100 12,077 12,577
A0A2Z5U3Z0_9INFA_Doud_2016 [doud2016accurate] 400 100 10,215 10,715
A0A2Z5U3Z0_9INFA_Wu_2014 [wu2014high] 400 100 1,850 2,350
A4D664_9INFA_Soh_2019 [soh2019comprehensive] 400 100 13,921 14,421
C6KNH7_9INFA_Lee_2018 [lee2018deep] 400 100 10,254 10,754
CAPSD_AAV2S_Sinai_2021 [sinai2021generative] 400 100 41,828 42,328
I6TAH8_I68A0_Doud_2015 [doud2015site] 400 100 8,962 9,462
PA_I34A1_Wu_2015 [wu2015functional] 400 100 1,320 1,820
POLG_DEN26_Suphatrakul_2023 [suphatrakul2023functional] 400 100 16,397 16,897
Q2N0S5_9HIV1_Haddox_2018 [haddox2018mapping] 400 100 12,229 12,729
R1AB_SARS2_Flynn_2022 [flynn2022comprehensive] 400 100 5,225 5,725
RDRP_I33A0_Li_2023 [li2023deep] 400 100 11,503 12,003
REV_HV1H2_Fernandes_2016 [fernandes2016functional] 400 100 1,647 2,147
TAT_HV1BR_Fernandes_2016 [fernandes2016functional] 400 100 1,077 1,577